One computer can process the payroll for a company in 5hours. A new computer can process it in 3 hours. Which equation could be used to find how long it will take to process the payroll if both computers are used.
Answer by nerdybill(7384)

One computer can process the payroll for a company in 5hours. A new computer can process it in 3 hours. Which equation could be used to find how long it will take to process the payroll if both computers are used.
.
Since you don't provide the options available, the general equation is:
Let x = time (hours) for both
then
x(1/5 + 1/3) = 1
to solve:
multiplying both sides by 15:
x(3 + 5) = 15
x = 15/8
x = 1 hour and 53 minutes